RikeyahF:

Simplify:(2x+8)(7x^2-9x-5)

RikeyahF:

A 14x^3+74x^2+62x-40 B 14x^3+38x^2+62x-40 C 14x^3+38x^2+82x-40 D 14x^3+74x^2+82x-40

umm:

Starting with this\[(2x+8)(7x ^{2-}9x-5)\] \[=(2x+8)(7x ^{2}+−9x+−5)\]\[=(2x)(7x ^{2})+(2x)(−9x)+(2x)(−5)+(8)(7x ^{2})+(8)(−9x)+(8)(−5)\]\[=14x ^{3}−18x ^{2}−10x+56x ^{2}−72x−40\]\[=14x ^{3}+38x ^{2}−82x−40\] Final answer\[=14x ^{3}+38x ^{2}−82x−40\]

RikeyahF:

Thank you🤍

umm:

Anytime. Welcome to QuestionCove, by the way.

RikeyahF:

Thank you 🙏🏾

umm:

(: